Two 2.5-cm-diameter-disks spaced 1.5 mm apart form a parallel-plate capacitor. The electric field between the disks is 4.5×105 V/m .

An electron is launched from the negative plate. It strikes the positive plate at a speed of 2.5×107 m/s . What was the electron's speed as it left the negative plate?
Answer #1

V = E*d = 4.5*10^5 * 1.5*10^-3 =675 Volt

By energy conservation suppose electron speed as it leaves the -ve plate =vi

1/2*m*vi^2 = qΔV + 1/2*m*vf^2

charge of the electron q = -1.6*10^-19 C ,

and its mass m = 9.11*10^-31 kg

0.5*9.11*10^-31*vi^2 = -1.6*10^-19*675 + 1/2*9.11*10^-31* ( 2.5*10^7)^2

v≈1.96951×10^7 m/s answer  

Goodluck for exams Comment in case any doubt, will reply for sure..